# Gaussian Scenes: Pose-Free Sparse-View Scene Reconstruction using Depth-Enhanced Diffusion Priors

### v1

@@ 3.3 Pixel-Aligned Confidence Map
To effectively guide our diffusion model towards recognizing and repairing artifacts in novel views, we design a novel pixel-aligned confidence measure based on the cumulative transmittance along a viewing direction at a pixel $i$ and the number of overlapping Gaussians at that pixel. Formally, we define confidence at a pixel as:
$\mathcal{C}_{i}=-\log(T_{i}+\epsilon)\times n_{contrib}$ (4)
where $T_{i}=\prod_{i}(1-\alpha_{i})$ represents the fraction of light transmitted through pixel $i$ without interaction and $n_{contrib}$ is the number of contributing 3D Gaussians. $\epsilon>0$ is a small constant to prevent taking the logarithm of 0. Intuitively, low $T_{i}$ means the viewing ray interacts significantly with 3D Gaussians, implying higher confidence in the pixel’s color. The negative logarithm provides a more perceptually uniform confidence scale. $n_{contrib}$ further increases (or decreases) the reliability of the rendered color or depth due to consensus among multiple (few) Gaussian primitives.
To the best of our knowledge, only 3DGS-Enhancer (Liu et al., 2024) proposes a confidence measure based on alpha-compositing of Gaussian scales, assuming that well-reconstructed areas are represented by Gaussians of small volumes, whereas large scales usually point towards artifacts. However, as can be seen in Fig 2, this assumption does not hold true in the case of regions with monotonous textures, which do not necessarily need multiple Gaussians of small scales to represent fine details. One can clearly see that our proposed measure effectively captures both empty regions as well as reconstruction artifacts, avoiding false positives unlike  Liu et al. (2024).
Render Conf. Measure  (Liu et al., 2024) Conf. Measure (Ours)
DTU scan38 (6) Refer to caption Refer to caption Refer to caption
DTU scan25 (6) Refer to caption Refer to caption Refer to caption
@@
Figure 2: Confidence Measure comparison with 3DGS-Enhancer (Liu et al., 2024). Our confidence map accurately identifies artifacts and 00-Gaussian regions in the image (darker pixels) while  (Liu et al., 2024) incorrectly attributes high confidence to regions with overlap of small-scale Gaussians.
3.3.1 Synthetic RGBD Dataset Creation
Figure 3: Training our RGBD diffusion model. Pairs of RGBD clean images and images with artifacts are obtained from 3DGS fitted to sparse and dense observations, respectively, across $165$ scenes. CLIP features provide scene context, and a confidence map additionally detects empty regions and artifacts in the artifact image. These are combined with one of 51 synonymous prompts generated by GPT-4 Achiam et al. (2023) from a base instruction. The Stable Diffusion UNet is then fine-tuned with a dataset of $28,161$ samples.
@@ 3.3.1 Synthetic RGBD Dataset Creation
For training the additional weights in $\bm{\epsilon_{\theta}}$ for RGBD image-to-image diffusion, we rely on a set $\mathcal{X}=\{(I^{i},\hat{I}^{i},\mathcal{C}^{i},y^{i})_{i=1}^{N}\}$ of data quartets, each containing a clean RGBD image $I^{i}$, an RGBD image with artifacts $\hat{I}^{i}$ and the corresponding confidence map $\mathcal{C}^{i}$, and an edit instruction for fine-tuning $y^{i}$, to “teach” the diffusion model how to detect Gaussian artifacts guided by the confidence map and generate a clean version of the conditioning image. For this, we build an artifact simulation engine comprising a high-quality 3DGS model fitted to dense views, a low-quality 3DGS model fitted to few views, and camera interpolation and perturbation modules to use supervision of the high-quality model at viewpoints beyond ground truth camera poses. The fine-tuning setup is illustrated in Fig. 3. For a given scene, we fit sparse models for $M\in\{3,6,9,18\}$ number of views. To have diversity in edit instructions, we start with a base human-intuitive instruction - “Inpaint low-confidence regions in the image and depth map with coherent details and remove all floaters and Gaussian artifacts.” and ask GPT-4 to generate 50 synonymous instructions. During training, each image triplet is randomly combined with one of these 51 instructions. Additionally, for each $M$, we save the CLIP features $c_{clip}$ of the $M$ sparse views for conditioning $\bm{\epsilon_{\theta}}$
3.3.2 Depth-Augmented Autoencoder Finetuning
@@
$\mathcal{L}_{\text{autoencoder}}=$ $\underset{\mathcal{E},D}{\text{min}}\,\underset{D_{\psi}}{\text{max}}\big{(}\mathcal{L}_{\text{rec}}(x,D(\mathcal{E}(x)))$ (5)
$-\mathcal{L}_{\text{adv}}(D(\mathcal{E}(x)))$
$+\log(D_{\psi}(x))$
$+\mathcal{L}_{\text{reg}}(x;\mathcal{E},D)\big{)}$
@@ 3.3.3 UNet Finetuning
With our finetuned autoencoder, we next train the UNet with the frozen VAE on $\mathcal{X}$ with the following objective:
$\mathcal{L}=\mathbb{E}_{i\sim\mathcal{U}(N),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}\Big{[}$ $\|\epsilon_{t}-\bm{\epsilon_{\theta}}(\mathbf{z}^{i}_{t};t,\mathcal{E}(\hat{I}),\mathcal{\hat{C}},c_{attn},c_{\theta}(y^{i}))\|_{2}^{2}\Big{]}$ (6)
Input RGB Recon. RGB Input Depth Recon. Depth
where $\mathbf{z}_{t}^{i}$ is the encoded image ${I}^{i}$ diffused with sampled noise $\epsilon$ at time step $t$. Thus, the UNet learns to generate clean RGBD images $I^{i}$, conditioned on artifact RGBD images $I^{i}$, confidence maps $\mathcal{C}^{i}$, CLIP features $c_{attn}$, and instructions $y^{i}$.
3.3.4 Inference using RGBD Novel View Synthesis
At inference time, given a render and depth map with artifacts and the base instruction $y=$“Inpaint low-confidence regions in the image and depth map with coherent details and remove all floaters and Gaussian artifacts." the finetuned UNet $\bm{\epsilon_{\theta}}$ learns to predict the noise in latent $\mathbf{z}_{t}$ according to $t\sim\mathcal{U}[t_{min},t_{max}]$ as:
$\hat{\epsilon}_{t}=$ $\,\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\mathcal{\hat{C}},c_{attn},\varnothing)+s_{I}(\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\mathcal{E}(\hat{I}),\mathcal{\hat{C}},c_{attn},\varnothing)-\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\mathcal{\hat{C}},c_{attn},\varnothing))$ (7)
$+s_{T}(\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\mathcal{\hat{C}},c_{attn},c_{\theta}(y))-\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\mathcal{E}(\hat{I}),\mathcal{\hat{C}},c_{attn},\varnothing))$
where $s_{I}$ and $s_{T}$ are the RGBD image and prompt guidance scales, dictating how strongly the final multistep reconstruction agrees with the RGBD render $\hat{I}$ and the edit prompt $y$, respectively. After $k=20$ DDIM Song et al. (2021) sampling steps, we obtain our final RGBD render by decoding the denoised latent as $x_{\pi}=[I_{\pi},D_{\pi}]=\mathcal{D}(z_{0})$.
3.4 Scene Reconstruction with Diffusion Priors
Our diffusion priors infer plausible detail in unobserved regions. However, the generated images at novel poses lack 3D consistency due to lack of explicit 3D conditioning of the UNet through pose embeddings. For this, we devise an iterative strategy where we first sample novel views along a B-spline trajectory fitted to the training views. We initialize the Gaussian optimization with the set of Gaussians $\mathcal{G}$ fitted to the training views. Each novel view is added to the training stack at the beginning of every densification step to encourage the optimization to adjust to the distilled scene priors. At every iteration, we sample either an observed or unobserved viewpoint from the current training stack. We jointly optimize Gaussian attributes, and both train and novel poses. We employ the 3DGS objective for the training views. For novel views, we employ the SparseFusion Zhou and Tulsiani (2023) objective in the RGB space (Wu et al., 2024) and a PCC loss for the rendered and denoised depths.
@@ 3.4 Scene Reconstruction with Diffusion Priors
where $\mathcal{L}_{p}$ is the perceptual loss Zhang et al. (2018), $w(t)$ a noise-dependent weighting function, $I_{\pi},D_{\pi}$ are the rendered image and depth at novel viewpoint $\pi$, and $\hat{I}_{\pi},\hat{D}_{\pi}$ are their rectified versions obtained with our diffusion prior.
@@ 3.5 Test-time Pose Alignment
GScenes reconstructs a plausible 3D scene from unposed source images. However, reconstruction with few views is inherently ambiguous as several solutions can satisfy the train view constraints. Hence, the reconstructed scene would most likely be quite different from the actual scene from which $M$ views were sampled. Hence, for a given set of test views, following prior work Fan et al. (2024); Jiang et al. (2024), we freeze the Gaussian attributes and optimize the camera pose for each target view by minimizing a photometric loss between the rendered image and test view. Following this alignment step performed for $500$ iterations per test image, we evaluate the NVS quality.
InstantSplat RegNeRF DiffusioNeRF ZeroNVS* COGS Ours Ground Truth
@@
Treehill(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparison of GScenes with few-view methods. Our approach consistently fairs better in recovering image structure from foggy geometry, where baselines typically struggle with “floaters” and color artifacts.
@@ 4 Experiments
We compare GScenes with state-of-the-art pose-free and pose-required sparse-view reconstruction methods in Fig and Table . We also ablate the different components and design choices of our diffusion model.
@@ 4.1 Experimental Setup
We evaluate GScenes on the 9 scenes of the MipNeRF360 dataset (Barron et al., 2022), comprising 5 outdoor and 4 indoor scenes. Each scene has a central object or area of complex geometry with an equally intricate background. This makes it the most challenging $360^{\circ}$ dataset compared to CO3D (Reizenstein et al., 2021), RealEstate 10K (Zhou et al., 2018), DTU (Jensen et al., 2014), etc. We pick the $M$-view splits as proposed by ReconFusion Wu et al. (2024) and CAT3D Gao* et al. (2024) and evaluate all baselines on the official test views where every $8^{th}$ image is held out for testing.
@@
We fine-tune our diffusion model on a mix of 165 scenes encompassing Tanks and Temples (Knapitsch et al., 2017), CO3D (Reizenstein et al., 2021), DTU (Jensen et al., 2014), Deep Blending (Hedman et al., 2018), and LLFF (Mildenhall et al., 2019) to obtain a total of $28161$ data samples. We first train 3DGS on sparse and dense subsets of each scene for $M\in\{3,6,9,18\}$. For $M>18$, novel view renders and depth maps mostly show Gaussian blur as artifacts. Finetuning this model takes about $1$-day on a single A6000 GPU.
@@
We compare our approach against $8$ baselines. FreeNeRF (Yang et al., 2023), RegNeRF (Niemeyer et al., 2022), DiffusioNeRF (Wynn and Turmukhambetov, 2023) are pose-required few-view regularization methods based on NeRFs. ZeroNVS (Sargent et al., 2024) reconstructs a complete 3D scene from a single image using a novel camera normalization scheme and anchored SDS loss. We use the ZeroNVS baseline introduced in ReconFusion (Wu et al., 2024), designed to adapt ZeroNVS to multi-view inputs. ReconFusion and CAT3D are state-of-the-art multi-view conditioned diffusion models for sparse-view reconstruction. We use the reported average performance on the $9$ scenes of MipNeRF360 with classical metrics - PSNR, SSIM, and LPIPS for quantitative comparison and pick the relevant scenes and test views from the two papers for qualitative comparison. For pose-free methods, we pick InstantSplat (Fan et al., 2024) and COGS (Jiang et al., 2024) and show that our approach outperforms both in NVS qualitatively and quantitatively.

### v2

@@ 2.3 Pixel-Aligned Confidence Map
To guide our generative diffusion model in identifying problematic and artifact-ridden regions in novel views, we introduce a pixel-aligned confidence measure combining transmittance with Gaussian density:
$\mathcal{C}_{i}=-\log(T_{i}+\epsilon)\times n_{\text{contrib}}$ (2)
MASt3R + 3DGS Diffusion Sample (Conf. Map (Liu et al., 2024)) Diffusion Sample (Ours)
3 views Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 6: As an ablation for our confidence measure, we train a variant of our diffusion model using the confidence measure proposed in 3DGS Enhancer (Liu et al., 2024). Conditioning the UNet with this inaccurate confidence measure leads to implausible NVS.
where $T_{i}=\prod_{i}(1-\alpha_{i})$ represents light transmission without Gaussian interaction, $n_{\text{contrib}}$ counts contributing 3D Gaussians, and $\epsilon>0$ prevents logarithmic singularities. This formulation captures two complementary reliability signals: (1) low transmittance indicates significant Gaussian interactions, suggesting higher rendering confidence, and (2) consensus among multiple Gaussians validates pixel reliability through primitive agreement.
Unlike 3DGS-Enhancer (Liu et al., 2024), (one of the few methods that propose confidence measures for sparse-view scene reconstruction), which assumes that well-reconstructed areas contain small-scale Gaussians, our measure remains effective for monotonous textures, where fine-grained Gaussian representation is unnecessary. Such regions are usually represented by a few high-opacity large-scale Gaussians (low $n_{contrib}$, but also low $T_{i}$) and hence not flagged as low-confidence regions by our confidence measure. Fig. 7 demonstrates how our approach accurately identifies both empty regions and reconstruction artifacts while avoiding false positives. The significance of this improved confidence measure is evident in Fig. 6, where models trained with previous confidence formulations produce implausible novel views due to misleading confidence signals.
Sparse Render Conf. Measure  (Liu et al., 2024) Conf. Measure (Ours) Dense Render
3 views Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 7: Confidence Measure comparison with 3DGS-Enhancer (Liu et al., 2024). Our confidence map accurately identifies artifacts and 00-Gaussian regions in the sparse-view (darker pixels) while  Liu et al. (2024) incorrectly attributes high confidence to regions with overlap of small-scale Gaussians. NVS render from a densely fitted 3DGS representation is provided for reference.
Input RGB Recon. RGB Input Depth Recon. Depth
2.4 RGBD Dataset Creation
@@
For training the additional weights in $\bm{\epsilon_{\theta}}$ for RGBD image-to-image diffusion, we rely on a set $\mathcal{X}=\{(I^{i},\hat{I}^{i},\mathcal{C}^{i},c_{clip}^{i},c_{geo}^{i})_{i=1}^{N}\}$, each containing a clean RGBD image $I^{i}$, an RGBD image with artifacts $\hat{I}^{i}$ and the corresponding confidence map $\mathcal{C}^{i}$, CLIP features of source images $c_{clip}^{i}\in\mathcal{R}^{M\times 768}$, and plücker embeddings of source and target cameras $c_{geo}^{i}\in\mathcal{R}^{(M+1)\times 78}$, to “teach” the diffusion model how to inpaint missing details and detect Gaussian artifacts guided by the confidence map, context and geometry features and generate a clean version of the conditioning image. For this, we build a dataset generation pipeline comprising a high-quality 3DGS model fitted to dense views, a low-quality 3DGS model fitted to few views, and camera interpolation and perturbation modules to use supervision of the high-quality model at viewpoints beyond ground truth camera poses. For a given scene, we fit sparse models for $M\in\{3,6,9,18\}$ number of views. We render $I^{i}$ using the high-quality model and $\hat{I}^{i}$, $\mathcal{C}^{i}$ using the low-quality model. We save the CLIP features $c_{clip}$ and plücker embeddings of the $M$ sparse views and 1 target view per sample for conditioning $\bm{\epsilon_{\theta}}$ Our fine-tuning setup is illustrated in Fig  12 (Sec C).
2.5 Depth-Augmented Autoencoder Finetuning
@@
$\mathcal{L}_{\text{autoencoder}}=$ $\underset{\mathcal{E},D}{\text{min}}\,\underset{D_{\psi}}{\text{max}}\big{(}\mathcal{L}_{\text{rec}}(x,D(\mathcal{E}(x)))-\mathcal{L}_{\text{adv}}(D(\mathcal{E}(x)))+\log(D_{\psi}(x))+\mathcal{L}_{\text{reg}}(x;\mathcal{E},D)\big{)}$ (3)
@@ 2.5.1 UNet Finetuning
With our finetuned autoencoder, we next train the UNet with the frozen VAE on $\mathcal{X}$ with this objective:
$\mathcal{L}_{diff}=\mathbb{E}_{i\sim\mathcal{U}(N),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}\Big{[}$ $\|\epsilon_{t}-\bm{\epsilon_{\theta}}(\mathbf{z}^{i}_{t};t,\mathcal{E}(\hat{I}),\mathcal{\hat{C}},\mathcal{E}(\hat{I}\cdot\mathcal{C}),c^{clip}_{attn},c^{geo}_{attn})\|_{2}^{2}\Big{]}$ (4)
2.5.2 RGBD Novel View Synthesis
At inference time, given a render and depth map with artifacts, and confidence map, CLIP features and camera embeddings for conditioning, the finetuned UNet $\bm{\epsilon_{\theta}}$ learns to predict the noise in latent $\mathbf{z}_{t}$ according to $t\sim\mathcal{U}[t_{min},t_{max}]$ as:
$\hat{\epsilon}_{t}=$ $\,\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\varnothing,\varnothing,c^{clip}_{attn},c^{geo}_{attn})$ (5)
$+s_{I}(\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\mathcal{E}(\hat{I}),\mathcal{\hat{C}},\mathcal{E}(\hat{I}\cdot\mathcal{C}),c^{clip}_{attn},c^{geo}_{attn})-\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\mathcal{\hat{C}},\varnothing,c^{clip}_{attn},c^{geo}_{attn}))$
$+s_{C}(\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\mathcal{\hat{C}},\varnothing,c^{clip}_{attn},c^{geo}_{attn})-\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\varnothing,\varnothing,c^{clip}_{attn},c^{geo}_{attn}))$
where $s_{I}$ and $s_{C}$ are the RGBD image and confidence map guidance scales, dictating how strongly the final multistep reconstruction agrees with the RGBD render $\hat{I}$ and the confidence map $\mathcal{C}$, respectively. After $k=20$ DDIM (Song et al., 2021) sampling steps, we obtain our final RGBD render by decoding the denoised latent as $x_{\pi}=[I_{\pi},D_{\pi}]=\mathcal{D}(z_{0})$.
2.6 Scene Reconstruction with Generative Priors
Our generative diffusion model provides a generative prior to infer plausible detail in unobserved regions. Despite view conditioning using pose embeddings, the generated images at novel poses lack complete 3D consistency. For this, we devise an iterative strategy where we first sample novel views along an elliptical trajectory fitted to the training views. We initialize the Gaussian optimization with the set of Gaussians $\mathcal{G}$ fitted to the training views (Sec 2.1). Each novel view is added to the training stack at the beginning of every densification step to encourage the optimization to adjust to the distilled scene priors. At every iteration, we sample either an observed or unobserved viewpoint from the current training stack. We bring back Adaptive Density Control to encourage densification of Gaussians in 0-Gaussian regions. We employ the 3DGS objective for the training views. For novel views, we employ the SparseFusion (Zhou & Tulsiani, 2023) objective in the RGB space and a PCC loss for the rendered and denoised depths.
@@ 2.6 Scene Reconstruction with Generative Priors
where $\mathcal{L}_{p}$ is the perceptual loss (Zhang et al., 2018), $w(t)$ a noise-dependent weighting function, $I_{\pi},D_{\pi}$ are the rendered image and depth at novel viewpoint $\pi$, and $\hat{I}_{\pi},\hat{D}_{\pi}$ are their rectified versions obtained with our diffusion prior. The PCC loss is defined as $PCC(D_{\pi},\hat{D}_{\pi})=1-\frac{\operatorname{Cov}(D_{\pi},\hat{D}_{\pi})}{\sigma_{D_{\pi}}\sigma_{\hat{D}_{\pi}}}$.
@@ 2.7 Test-time Pose Alignment
GScenes reconstructs a plausible 3D scene from pose-free source images. However, reconstruction with few views is inherently ambiguous as several solutions can satisfy the train view constraints. Hence, the reconstructed scene would probably be quite different from the actual scene from which $M$ views were sampled. Hence, for a given set of test views, following prior work (Fan et al., 2024; Jiang et al., 2024), we freeze the Gaussian attributes and optimize the camera pose for each target view by minimizing a photometric loss between the rendered image and test view. Following this alignment step performed for $500$ iterations per test image, we evaluate the NVS quality.
[…]
Bonsai(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Counter(3) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 9: Qualitative comparison of GScenes with few-view methods on the MipNeRF360 dataset. Our approach consistently fairs better in recovering image structure from foggy geometry, where baselines typically struggle with “floaters” and color artifacts.
[…]
3 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
6 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Qualitative comparison of GScenes with few-view methods on the DL3DV benchmark. Our method achieves plausible reconstruction in unobserved areas of complex scenes where even posed reconstruction techniques struggle.
@@ 3 Experiments
We compare GScenes with state-of-the-art pose-free and posed sparse-view reconstruction methods in Fig 9, 10 and Table 1, 2. We also ablate the different components and design choices of our diffusion model.
@@ 3.1 Experimental Setup
We evaluate GScenes on the 9 scenes of the MipNeRF360 dataset (Barron et al., 2022), and 15 scenes (out of 140) of the DL3DV-10K benchmark dataset. For MipNeRF360, We pick the $M$-view splits as proposed by ReconFusion and CAT3D and evaluate all baselines on the official test views where every $8^{th}$ image is held out for testing. For DL3DV-10K scenes, we create $M$-view splits using a greedy view-selection heuristic for maximizing scene coverage given a set of dense training views, similar to the heuristic proposed in Wu et al. (2024). For test views, we hold out every $8^{th}$ image as in MipNeRF360. Additionally, we pick the plant scene of CO3D for qualitative comparison with ReconFusion and CAT3D.
@@
We fine-tune our diffusion model on a mix of 1043 scenes encompassing Tanks and Temples (Knapitsch et al., 2017), CO3D (Reizenstein et al., 2021), Deep Blending (Hedman et al., 2018), and the $1k$ subset of DL3DV-10K (Ling et al., 2024) to obtain a total of $171,461$ data samples. We first train 3DGS on sparse and dense subsets of each scene for $M\in\{3,6,9,18\}$. For $M>18$, novel view renders and depth maps mostly show Gaussian blur as artifacts. Finetuning this model takes about $4$-days on a single A6000 GPU.
@@
We compare our approach against $8$ baselines, but our main comparison is with 2 pose-free methods - COGS and our 3D reconstruction engine - MASt3R + 3DGS. The other 6 methods are posed and included for a stronger comparison. Out of these, FreeNeRF, RegNeRF, and DiffusioNeRF are sparse-view regularization methods based on NeRFs. ZeroNVS reconstructs a complete 3D scene from a single image using a novel camera normalization scheme and anchored SDS loss. We use the ZeroNVS baseline introduced in ReconFusion, designed to adapt ZeroNVS to multi-view inputs. Further, we also provide the reported quantitative results from ReconFusion and CAT3D, which are closed source and unverifiable, to show that our method gets close to their performance despite using a fraction of their compute. We also pick the relevant scenes and test views from the 2 papers for qualitative comparison.
Render Base Base + Conf. w/o CLIP
Bonsai(3) Refer to caption Refer to caption Refer to caption Refer to caption
w/o Conf. Cross-Attn Ours Ground Truth
Bonsai(3) Refer to caption Refer to caption Refer to caption Refer to caption
Render Base Base + Conf. w/o CLIP
Treehill(3) Refer to caption Refer to caption Refer to caption Refer to caption
w/o Conf. Cross-Attn Ours Ground Truth
Treehill(3) Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Ablation Study with the 3-view splits of Bonsai and Treehill scenes from MipNeRF360. Images show samples from different variants of our diffusion model based on conditioning.